OpenStudy (anonymous):

Jules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again?

OpenStudy (anonymous):

depends on the angle at which the ball is hit in the air

OpenStudy (anonymous):

v=gt t=v/g=2.5 s=1/2*2.5*2=3.125feet

myininaya (myininaya):

s(t)=-16t^2+25t+0 is the formula for height